how do you simplify raticals that have an equationon both sides like
X-3=THE SQUARE ROOT OF 30-2X
AND CAN YOU GIVE ME SOME OTHER EXAMPALES.
THANK YOU,
FROM,
MATH BREAKDOWN

Answer by checkley71(8403)

SQUARE BOTH SIDES OF THE EQUATION THUS X~2-6X+9=30-2X OR X~2-4X-21 OR
(X-7)(X+3)=0 OR X=7 & X=-3